OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (anonymous):

y-y1 = M(x-x1) m is the slope plug the numbers into the equation

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(1\quad ,&-4) \end{array} \\\quad \\ slope = m= \cfrac{1}{2} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)